OpenStudy (anonymous):

Divide eight and one-sixth divided by four and two-thirds. two and one-half four-sevenths two and two-thirds one and three-fourths

OpenStudy (imstuck):

\[8\frac{ 1 }{6 }\div 4\frac{ 2 }{3 }->\frac{ 49 }{ 6 }\times \frac{ 3 }{ 14 }\]you see what I did there?

OpenStudy (imstuck):

change them to improper fractions and then change the sign and flip the second fraction in order to multiply.

OpenStudy (anonymous):

@mkerr017

OpenStudy (anonymous):

answer plz

OpenStudy (imstuck):

one and three-fourths.

OpenStudy (anonymous):

thx

OpenStudy (imstuck):

You're welcome.

OpenStudy (anonymous):

sorry, i was eating dinner:/